# Audio Captioning via Generative Pair-to-Pair Retrieval with Refined Knowledge Base

### v1

@@ 1 Introduction
In image captioning, the RAG approach has been widely adopted to take advantage of external knowledge and improve performance (Ramos et al. 2023; Li et al. 2024; Ramos, Elliott, and Martins 2023). Although recent studies have begun to apply RAG to audio captioning (Ghosh et al. 2024; Kong et al. 2024), there has been limited exploration of how to effectively adapt this approach to the specific challenges of the audio captioning. One of the key challenges when applying RAG to audio captioning is the inherent ambiguity of audio signals (Martin Morato and Mesaros 2021; Lipping, Drossos, and Virtanen 2019; Drossos, Lipping, and Virtanen 2020), where the same audio can be interpreted in multiple ways depending on the context and the listener’s perception. This ambiguity complicates the task of retrieving relevant information that can enhance caption generation, as the retrieved data may not always align with the contextualized intent of the audio query. For example, the audio caption “a person is talking over an intercom as vehicles and a train pass by in the background” could be described as “background noise is present” or “faint talking is at an event with music in the background.” Apparently, the same audio can be described differently as its intent may vary depending on the context and situation in which it is used. If the caption of a retrieved audio is described as “background noise is present,” it may not provide the necessary context for accurate generation. The fact that audio can be described differently can reduce the consistency of audio captioning. Moreover, this issue can lead to performance degradation in real-world scenarios where the contextualized intent differs drastically, such as distinguishing between a construction site and a hospital. Previous methods have typically relied solely on unimodal audio queries, using simple retrieval strategies and a basic knowledge base.
However, these approaches, as mentioned earlier, do not take into account the diverse interpretations that can arise from the same audio input. Therefore, it is necessary to explore how the RAG approach can be adapted to address the ambiguity of audio data, ensuring that the retrieved information considers not only the audio modality but also the text modality for generating captions that accurately reflect the contextualized intent of the audio query.
@@
To address these challenges, we propose a generative pair-to-pair retrieval method combined with a refined knowledge base for audio captioning. Unlike traditional unimodal queries, the pair-to-pair retrieval method utilizes multi-modal queries, taking both modalities into account. To determine whether considering both modalities simultaneously has a direct impact on performance, we analyze the similarity between audio and text using various retrieval methods and knowledge bases. Our findings demonstrate that audio similarity alone is insufficient for accurately retrieving audio-text pairs. Our generative pair-to-pair retrieval method transforms audio queries into text queries and uses both modalities as queries to retrieve audio-text pairs. Additionally, our refined knowledge base is constructed with candidate pairs that consider both modalities and align with contextualized intents. Using these two approaches, we address the ambiguity in the audio data and enhance the overall performance of the audio captioning.
@@
  • Second, to retrieve these beneficial audio-text pairs, we propose a generative retrieval method and a refined knowledge base, both utilizing pair-to-pair search.
@@ 3 Method
In this section, we propose an effective RAG-based approach for audio captioning. We describe the model architecture and training methodology, and introduce two key methods: generative pair-to-pair retrieval and a refined knowledge base, both of which utilize audio-text query to use pair-to-pair retrieval.
@@ 3.1 Model Architecture
The model architecture consists of three main components: a pre-trained audio encoder, a language model for text generation, and a projection network to map audio features into the language model’s embedding space. When an audio query is received, the Laion-CLAP (Wu et al. 2023) encoder extracts features and maps them into the language model embedding space via an MLP projection layer as a connector.
The audio encoder processes variable-length audio input, converts them into mel-spectrogram features with a 1024 window length, and projects the audio features to the language model as audio tokens. To reduce computational complexity, we decrease the number of audio tokens by applying average pooling with a stride, balancing important audio details and efficiency. For audio inputs longer than 30 seconds, a special length token is added before passing the input to the Vicuna (Chiang et al. 2023) language model, which then generates the caption. The model also retrieves relevant audio-text pairs to provide contextual information, improving the accuracy of the caption. Details on the optimal number of audio tokens are provided in the Appendix A.
@@ 3.2 Training Methods
We adopt two-stage training methods following recent approaches (Li et al. 2022, 2023; Liu et al. 2023). In the first stage, we focus on aligning the audio modality with the language model. This involves training audio connector to map audio features into the language model embedding space, enabling the language model to interpret and generate accurate captions.
@@ 3.3 Pair-to-Pair Retrieval
We propose a pair-to-pair search method that uses a multimodal query, combining both an audio query and the corresponding text query, to find relevant pairs. This approach allows for the retrieval of pairs more accurately similar to the audio query by effectively utilizing both modalities. The corresponding text query contains contextualized intent related to the task, defining what should be extracted from the audio query. For example, in audio captioning, the text query describes key sound events in the audio.
@@
This approach ensures that both audio and text information is effectively leveraged to enhance the relevance and accuracy of the retrieved pairs, leading to improved performance in audio captioning tasks.
However, in audio captioning, where the input is audio and the text is the target of generation, utilizing the text query is not straightforward because only an audio is provided as the input. To address this problem, we propose a generative pair-to-pair approach that incorporates generated caption in the retrieval process, enabling the use of a multimodal query combining both audio and text. Additionally, we leverage pair-to-pair retrieval to construct a refined knowledge base by eliminating irrelevant candidates and extracting only those that are relevant.
@@ 3.4 Generative Pair-to-Pair Retrieval
To accurately retrieve relevant audio-text pairs for a given audio query, we propose Generative Pair-to-Pair retrieval for audio-text pairs considering both modalities. In this approach, the generated caption serves as an text query $T_{q}$, transforming the audio query into a text query which describes query audio content. For a given query audio $A_{q}$ and the generated caption $T_{q}$, we retrieve relevant pairs $(A_{k},T_{k})$ from the knowledge base using combined similarity scores of both modalities, as defined in Equation (4).
We hypothesize that generated caption can partially contain description of query audio. Our retrieval method aims to identify audio-text pairs with descriptions that closely align with the contextualized intent of the query audio. We formalize the generative query transformation process where an audio language model $f_{\theta}$ generates a text query $T_{q}$ from the given audio query $A_{q}$.
@@ 3.5 Refined Knowledge Base
Pair-to-Pair Retrieval:
@@
where $N$ is the total number of pairs in the knowledge base.
Filtered Knowledge Base:
The retrieved pairs are then used to construct the refined knowledge base M, ensuring that the audio-text pairs are aligned with the trainset. This filtering process enhances the relevance of the retrieval results.
We propose a pair-to-pair retrieval method that utilizes a multimodal query, considering both modalities, in contrast to traditional approaches that rely solely on unimodal queries. This method introduces generative pair-to-pair retrieval and a refined knowledge base. generative pair-to-pair retrieval aids in retrieving relevant pairs using generated captions as a text query, while the refined knowledge base extracts useful pairs to form a more accurate knowledge base. Both approaches effectively retrieve pairs that enhance audio captioning performance.
@@ 4.1 Datasets and Metrics
Knowledge Base: We utilize the FAISS (Johnson, Douze, and Jégou 2019) library to accelerate dense retrieval using Laion-CLAP to find similar audio with captions. For our experiments, we define four different knowledge base configurations. The base KB includes only the training dataset. The large KB extends the base KB by incorporating WavCaps, a large-scale dataset of synthetic captions. Furthermore, the huge KB builds upon the large KB by including AutoACD as total 2.2M pairs. Additionally, we introduce the refined KB with 127K pairs, which further refines the huge KB by aligning audio and text data more closely to high-quality train datasets. which closely aligns audio and text data with high-quality training datasets.
@@
Experiment Setup: We train model on stage 1 for 5 epochs with a learning rate of 1e-4 and a batch size of 256, using the AdamW optimizer and a cosine learning rate scheduler. For fine-tuning the AudioCaps and Clotho datasets, we trained for 10 epochs with a learning rate of 1e-5 and a batch size of 128, using the AdamW optimizer and a cosine learning rate scheduler. During training, we set the maximum number of retrieved samples to 5, and during testing, we evaluated up to 10 samples. We also fine-tuned the language model head with PEFT parameters while freezing the input embedding. In addition, we unfrozen the audio encoder for further training. We scaled the learning rate to 0.1 times for the audio encoder and the language model head. For curriculum learning, each phase was trained for 10 epochs with the same learning rate of 1e-5 and batch size of 128.
@@ 6.1 Knowledge Base
In Table 5, we analyze how refining the knowledge base impacts audio captioning performance by selecting candidate pairs closely aligned with high-quality training datasets. The results demonstrate that refining the knowledge base, even with a smaller version, consistently improves performance across different retrieval methods. For instance, in the AudioCaps dataset, the refined large KB outperforms the large KB in all retrieval scenarios. Interestingly, in the huge KB setting, performance decreased despite the increase in knowledge base size. This can be interpreted as follows: a larger knowledge base increases the number of candidate pairs available for retrieval, raising the possibility of retrieving relevant information for more accurate generation. However, adding noisy candidates can introduce irrelevant context, which may degrade performance. This experiment indicates that expanding the knowledge base size alone does not guarantee improved results.
@@ 6.3 Retreival Strategies
Examining Table 7 demonstrates that each audio and text similarity plays a critical role in RAG performance. In terms of audio similarity, a larger knowledge base more effectively retrieves audio-text pairs that closely align with the query audio, generally leading to improved overall performance. The results with a larger knowledge base outperform those with smaller ones. In particular, the text similarity of the retrieved pairs, which closely matches the query audio captions, generally provides stronger performance than audio similarity alone. The pair-to-pair retrieval method, which uses a generated caption as the text query, achieves the highest performance by balancing strong text similarity with good audio similarity, demonstrating its effectiveness on both the AudioCaps and Clotho datasets.

### v2

# Audio Captioning RAG via Generative Pair-to-Pair Retrieval with Refined Knowledge Base

@@ 1 Introduction
Although recent studies have begun to apply RAG to audio captioning (Ghosh et al. 2024; Kong et al. 2024), there has been limited exploration of how to effectively adapt this approach to the specific challenges of the audio captioning. One of the key challenges when applying RAG to audio captioning is the inherent ambiguity of audio signals (Martin Morato and Mesaros 2021; Lipping, Drossos, and Virtanen 2019; Drossos, Lipping, and Virtanen 2020), where the same audio can be interpreted in multiple ways depending on the context and the listener’s perception. This ambiguity complicates the task of retrieving relevant information that can enhance caption generation, as the retrieved data may not always align with the contextualized intent of the audio query.
For example, the audio caption “a person is talking over an intercom as vehicles and a train pass by in the background” could be described as “background noise is present” or “faint talking is at an event with music in the background.” Apparently, the same audio can be described differently as its intent may vary depending on the context and situation in which it is used. Previous methods have typically relied solely on unimodal audio query. these approaches do not take into account the diverse interpretations that can arise from the same audio input. For example, when an audio-to-audio approach retrieves ”background noise is present,” it may fail to provide the necessary context for accurate generation. The fact that audio can be described differently can reduce the consistency of audio captioning. Moreover, this issue can lead to performance degradation in real-world scenarios where the contextualized intent differs drastically, such as distinguishing between a construction site and a hospital. Therefore, it is necessary to explore how the RAG approach can be adapted to address the ambiguity of audio data, ensuring that the retrieved information considers not only the audio modality but also the text modality for generating captions that accurately reflect the contextualized intent of the audio query.
@@
To address these challenges, we propose P2PCAP, a retrieval-augmented LLM-based audio captioning method that resolves audio ambiguity by finding relevant audio-text pairs through multimodal queries. Unlike traditional approaches that rely on unimodal queries, P2PCAP uses pair-to-pair retrieval, combining audio and text modalities for improved accuracy. It employs generative pair-to-pair retrieval, transforming audio queries into text queries and using both audio and text as a combined query to retrieve audio-text pairs. Additionally, the refined knowledge base is constructed from candidate pairs that consider both modalities and align with contextualized intents. P2PCAP builds on a typical multimodal LLM architecture and training strategies to establish a robust baseline for evaluating retrieval performance within the RAG framework. It incorporates curriculum learning during a second-stage training process with interleaved audio-text pairs as input to enhance in-context few-shot capabilities. To assess the importance of considering both modalities, we analyze the similarity between audio and text using various retrieval methods and knowledge bases. Our results show that relying only on audio similarity is insufficient due to the inherent ambiguity of audio data, underscoring the need for multimodal retrieval strategies.
@@
  • Second, to retrieve these beneficial audio-text pairs, we propose a generative retrieval method and a refined knowledge base, both utilizing pair-to-pair retrieval.
@@ 3 Method
In this section, we present a RAG-based approach for audio captioning, focusing on the model design, training strategies, and retrieval methods with refined KB. Section 3.1 describes the model architecture, including the audio encoder, projection network, and language model. Section 3.2 outlines the two-stage training process to align audio modality and fine-tune the model with interleaved audio-text pair. Section 3.3 introduces pair-to-pair retrieval, leveraging multimodal queries to retrieve relevant audio-text pairs. Section 3.4 extends this to generative pair-to-pair retrieval, incorporating generated captions as text queries to enhance retrieval relevance. Finally, Section 3.5 presents the refined knowledge base, which filters irrelevant candidates to create a compact and effective KB for improving retrieval-augmented audio captioning.
Figure 1: P2PCAP architecture with two training stages: Stage 1 focuses on aligning audio and captions through fine-tuning projection network. Stage 2 incorporates interleaved audio-text inputs, aiming to enhance few-shot audio-captioning.
@@ 3.1 Model Architecture
As illustrated in Figure 1, our model architecture consists of three main components: a audio encoder, a language model for text generation, and a projection network to map audio features into the audio token that LLM can process. When an audio query is received, the Laion-CLAP (Wu et al. 2023) encoder extracts features and maps them into the language model embedding space via an MLP projection layer. When a query audio is provided, the audio encoder extracts and process mel-spectrogram features using a 1024 window length. These features are then passed to the projection network, which maps them into audio tokens for the language model. To reduce computational complexity, we decrease the number of audio tokens by applying average pooling with a stride, balancing important audio details and efficiency. We used a pooling size of 16 and a stride of 8 to reduce the audio tokens to 127. Appendix A provides details on the optimal number of tokens from ablation studies. For audio inputs longer than 30 seconds, a special length token is added to input of language model. The model processes interleaved audio-text pairs as input, consisting of relevant audio and corresponding text captions. These interleaved inputs are passed to the language model, which uses the combined context to generate a caption for the query audio. We utilize pre-trained networks, including Laion-CLAP (Wu et al. 2023) for audio encoder and Vicuna (Chiang et al. 2023) for language model.
@@ 3.2 Training Methods
We adopt two-stage training methods following recent approaches (Li et al. 2022, 2023; Liu et al. 2023). In the first stage, we align the audio modality with the language model by training the projection network to map audio features into audio tokens that the language model can process to generate accurate captions.
@@ 3.3 Pair-to-Pair Retrieval
We propose a pair-to-pair retrieval method that utilizes a multimodal query, considering both modalities, in contrast to traditional approaches that rely solely on unimodal queries. This method employs generative pair-to-pair retrieval and a refined knowledge base. Our generative pair-to-pair retrieval utilizes generated captions as text queries to enhance pair relevance, while the refined knowledge base selectively retains relevant audio-text pairs to construct a more compact and accurate knowledge base. Both approaches effectively retrieve pairs that enhance audio captioning performance.
As shown in Figure 2, we use the generated caption as the text query and construct a multimodal query using the generative pair-to-pair approach, following Equations (5) and (6). KB is constructed by concatenating these embeddings, as described in Equation (8), and subsequently refined using the method outlined in Equation (9). Finally, we retrieve relevant audio-text pairs by applying Equation (4), ensuring that the retrieved pairs effectively address the input audio query.
Figure 2: Our retrieval method is composed of two component: 1. Generative Pair-to-Pair Retrieval: The audio language model generates a caption $T_{g}$ from the audio query $A_{q}$, forming a multimodal query using the query audio and generated caption. This query retrieves relevant audio-text pairs from the knowledge base. 2. Refined Knowledge Base: To create a compact and accurate knowledge base, we utilize pair-to-pair retrieval with audio-text pairs relevant to the contextualized intent of the query audio.
@@
This approach ensures that both audio and text information is effectively leveraged to enhance the relevance and accuracy of the retrieved pairs, leading to improved performance in audio captioning tasks. Building on this, we detail two key components of our approach: Generative Pair-to-Pair retrieval, which incorporates generated captions to create multimodal queries, and the Refined Knowledge Base, which optimizes retrieval by filtering irrelevant candidates.
@@ 3.4 Generative Pair-to-Pair Retrieval
We propose Generative Pair-to-Pair retrieval for audio-text pairs considering both modalities. In audio captioning, where the input is audio and the text is the target of generation, utilizing the text query is not straightforward because only an query audio is provided as the input.
Our pair-to-pair retrieval uses generated caption from an audio captioning model that describe the query audio context. In this approach, the generated caption serves as an text query $T_{q}$, transforming the audio query into a text query which describes query audio content. For a given query audio $A_{q}$ and the generated caption $T_{q}$, we retrieve relevant pairs $(A_{k},T_{k})$ from the knowledge base using combined similarity scores of both modalities, as defined in Equation (4). Our retrieval method aims to identify audio-text pairs with descriptions that closely align with the contextualized intent of the query audio.
We formalize the generative query transformation process where an audio captioning model $f_{\theta}$ generates a text query $T_{q}$ from the given audio query $A_{q}$.
@@ 3.5 Refined Knowledge Base
Refine Knowledge Base:
@@
where $N$ is the total number of pairs in the knowledge base. The retrieved pairs are then used to construct the refined knowledge base M, ensuring that the audio-text pairs are aligned with the trainset. This filtering process enhances the relevance of the retrieval results.
@@ 4.1 Datasets and Metrics
Knowledge Base: We utilize the FAISS (Johnson, Douze, and Jégou 2019) library to accelerate dense retrieval using Laion-CLAP to find similar audio with captions. For our experiments, we define four different knowledge base configurations. The base KB includes only the training dataset. The large KB extends the base KB by incorporating WavCaps, a large-scale dataset of synthetic captions. Furthermore, the huge KB builds upon the large KB by including AutoACD as total 2.2M pairs. Additionally, we introduce the refined KB with 127K pairs, which further refines the huge KB by aligning audio and text data more closely to contextualized intent. which closely aligns audio and text data with high-quality training datasets.
@@
Experiment Setup: We train the model on stage 1 for 5 epochs using a combination of WavCaps, AudioCaps, and Clotho datasets, with a learning rate of 1e-4 and a batch size of 256, utilizing the AdamW optimizer and a cosine learning rate scheduler. For fine-tuning, we exclusively use the AudioCaps and Clotho datasets, training for 10 epochs with a learning rate of 1e-5 and a batch size of 128, while maintaining the AdamW optimizer and cosine learning rate scheduler. During training, we set the maximum number of retrieved samples to 5, and during testing, we evaluated up to 10 samples. We also fine-tuned the language model head with PEFT parameters while freezing the input embedding. In addition, we unfrozen the audio encoder for further training. We scaled the learning rate to 0.1 times for the audio encoder and the language model head. For curriculum learning, each phase was trained for 10 epochs with the same learning rate of 1e-5 and batch size of 128.
@@ 6.1 Knowledge Base
In Table 5, we analyze how refining the knowledge base impacts audio captioning performance by selecting candidate pairs closely aligned with contextualized intent. The results demonstrate that refining the knowledge base, even with a smaller version, consistently improves performance across different retrieval methods. For instance, in the AudioCaps dataset, the refined large KB outperforms the large KB in all retrieval scenarios. Interestingly, in the huge KB setting, performance decreased despite the increase in knowledge base size. This can be interpreted as follows: a larger knowledge base increases the number of candidate pairs available for retrieval, raising the possibility of retrieving relevant information for more accurate generation. However, adding noisy candidates can introduce irrelevant context, which may degrade performance. This experiment indicates that expanding the knowledge base size alone does not guarantee improved results.
@@ 6.3 Retreival Strategies
Examining Table 7 demonstrates that each audio and text similarity plays a critical role in RAG performance. In terms of audio similarity, a larger knowledge base more effectively retrieves audio-text pairs that closely align with the query audio, generally leading to improved overall performance. The results with a larger knowledge base outperform those with smaller ones. In particular, the text similarity of the retrieved pairs, which closely matches the query audio captions, generally provides stronger performance than audio similarity alone. The generative pair-to-pair retrieval method, which uses a generated caption as the text query, achieves the highest performance by balancing strong text similarity with good audio similarity, demonstrating its effectiveness on both the AudioCaps and Clotho datasets.